find the values of k for which the following equations have no real roots.

(a) kx^2-4x+8=0

(b) 3x^2+5x+k+1=0

(c) 2x^2+8x-5=kx^2

(d) 2x^2+k=3(x-2)

(e) kx^2+2kx=4x-6

(f) kx^2+kx=3x-2


1
Expert's answer
2022-01-10T17:00:43-0500

(a)


"kx^2-4x+8=0"

"k\not=0"

"D=(-4)^2-4(k)(8)<0"

"16-32k<0"

"k>\dfrac{1}{2}"

"k\in(\dfrac{1}{2},\infin)"

(b)


"3x^2+5x+k+1=0"

"D=(5)^2-4(3)(k+1)<0"

"13-12k<0"

"k>\dfrac{13}{12}"

"k\in (\dfrac{13}{12}, \infin)"

(c)


"(2-k)x^2+8x-5=0"

"k\not=2"


"D=(8)^2-4(2-k)(-5)<0"

"16+10-5k<0"

"26-5k<0"

"k>\dfrac{26}{5}"

"k\in (\dfrac{26}{5}, \infin)"

(d)


"2x^2-3x+6+k=0"


"D=(-3)^2-4(2)(6+k)<0"

"9-48-8k<0"

"8k>-39"

"k>-\dfrac{39}{8}"

"k\in (-\dfrac{39}{8}, \infin)"



(e)


"kx^2+2kx-4x+6=0"

"k\not=0"


"D=(2k-4)^2-4(k)(6)<0"

"4k^2-16k+16-24k<0"

"k^2-10k+4<0"

"k^2-10k+25-21<0"

"(k-5)^2<21"

"5-\sqrt{21}<k<5+\sqrt{21}"

"k\in(5-\sqrt{21},5+\sqrt{21})"

(f)


"kx^2+kx=3x-2"

"k\not=0"


"kx^2+(k-3)x+2=0"

"D=(k-3)^2-4(k)(2)<0"

"k^2-6k+9-8k<0"

"k^2-14k+49<40"

"(k-7)^2<40"

"7-2\sqrt{10}<k<7+2\sqrt{10}"

"k\in (7-2\sqrt{10},7+2\sqrt{10})"